On the basis data from 1990 to 2006, the median income y in the year x for men and women is approximated by the equation given below, where x=0 corresponds to 1990 and y is in constant 2006 dollars. If these equations remain valid in the future in what year will the median income of men and women be the same?

Men: -245x+2y=58,990

Women: -847x+3y=48,484
Answer #1

Given Equations

Men: -245x+2y=58,990

Women: -847x+3y=48,484

Rearrange the equations in terms of median income (y)

2y = 58990 + 245x ...................(1)

3y = 48484 + 847x ...................(2)

simplifying equations we get

58990 + 245

y 29495 + 245 ....................(3)

48484847r..................(4)

Equating Equations (3) and (4)

245.r 48484 847r 29495245z

48484847r 245r 3

0001 1694r735.r 6

40001 959r 6

40001 959z

40001 * 2 = 959x

80002 = 959x

80002 959

x = 83.4 ~ 83 years

x=0 corresponds to 1990

So after 83 years that is 1990 + 83 = 2073

Hence the median income = 2073